OpenStudy (thanatos2154):

lim (sin(x))^(tan(x)) as x ->0+

OpenStudy (anonymous):

\[\large\begin{align*}\lim_{x\to0^+}(\sin x)^{\tan x}&=\lim_{x\to0^+}e^{\ln(\sin x)^{\tan x}}\\\\ &=\lim_{x\to0^+}e^{\tan x\ln(\sin x)}\\\\ &=\lim_{x\to0^+}e^{\frac{\ln(\sin x)}{\cot x}}\\\\ &=e^{\lim\limits_{~~x\to0^+}\frac{\ln(\sin x)}{\cot x}}\end{align*}\] The limit in the exponent is \(\dfrac{\infty}{\infty}\). Use L'Hopital's rule.

OpenStudy (thanatos2154):

Thx.